# Cavity modification of magnetoplasmon mode through coupling with intersubband polaritons

### v1

@@ I Introduction
In Landau polariton systems, the description of the coupling relies on Kohn’s theorem, which states that the cyclotron resonance frequency in a translationally-invariant two-dimensional gas is unaffected by electron-electron interactions [28]. Despite its simplicity, this result turns out to be extremely powerful, reducing the system to a non-interacting one, and allowing for the development of a simple, intuitive picture of light-matter interactions in terms of a Hopfield-like polaritonic description [29, 30]. At the same time, Kohn’s theorem is also a limitation, excluding non-linearities and possibly many-body physics from the optical properties of these systems. Going beyond its range of validity is thus becoming an important topic of research [31], which turns out to be surprisingly challenging: besides the use of intense THz sources as in Ref. [32], the fundamental strategy to break Kohn’s theorem and observe the effect of Coulomb interactions in the cyclotron resonance is to design the system to explicitly break translational invariance. This can be naturally achieved in resonant cavities with highly sub-wavelength field confinement and inhomogeneous fields [33], where the violation of Kohn’s theorem is manifested as a non-local polaritonic response [25, 34, 35].
@@
Using THz time-domain spectroscopy, we observe a pronounced renormalization of the magnetoplasmon resonance frequency through electron–electron Coulomb interactions that is activated by the spatial dependence of the polaritonic field arising from the hybridization of the TM cavity mode with the ISBT. Unlike Ref. [25], where non-locality resulted from a deeply sub-wavelength cavity miniaturization design ($\lambda/1000$), here the violation of Kohn’s theorem occurs with moderate field confinement ($\lambda/60$), provided that the resonant mode is sufficiently inhomogeneous. Furthermore, we demonstrate that by tuning the magnetoplasmon with the magnetic field, it is possible to strongly couple it with a different TE mode, which is mostly homogeneous and in-plane polarized. As a consequence, the TE mode is not hybridized with the ISBT, and thus exhibits purely cavity-like behavior. Moreover, the non-locality disappears and Kohn’s theorem is re-established due to a high level of spatial homogeneity.
@@
Figure 2: a) Schematic of experimental reflection THz-TDS set up. THz transmitter and receiver indicated with a ‘T’ and ‘R’ respectively. b) Measured reflectance spectra of the QW-loaded cavity as a function of temperature (with no applied magnetic field). The lower polariton branch (LP) is indicated. Below schematic indicating the frequency mismatch between the cavity and ISB resonances. Inset - single spectra slice at 10K. c) Measured THz-TDS transmission spectra of semiconductor heterostructure in the absence of a cavity at 3K. The black dashed line indicates the unperturbed MP frequency given by $\omega_{B}=eB/m^{*}$, where $m^{*}=0.067m_{e}$. Inset - single spectra slice at 3T.
@@ II.2 Quantum Well
A $3\mathrm{{\mu}m}$ stack of 53 GaAs/AlGaAs square quantum wells (red) with an electron density per well of $3\times 10^{11}\mathrm{cm^{-2}}$ is placed in the cavity below an $8\mathrm{{\mu}m}$ GaAs spacer. The QWs exhibit two resonances in the THz region, which we investigate in our study: the ISBT and the MP resonance.
@@
When a magnetic field is applied, the QW also exhibits an MP resonance due to the collective cyclotron resonance of the electrons [38]. Unlike the ISBT, which is fixed in frequency with the material growth, the MP resonance frequency increases linearly with applied magnetic field as $\omega_{B}=eB/m^{*}$, where $m^{*}=0.067m_{e}$ is the effective mass of electrons in the GaAs QW. The MP resonance measured in the bare QW heterostructure without the cavity is shown in Fig. 2c. The THz response shows an almost perfect agreement with the expected linear dependence on the magnetic field (shown as a dashed black line). The linewidth can be estimated to have an almost $B$-independent value of $\kappa/(2\pi)\approx 300\,$GHz (see inset), which is the expected order of magnitude considering the number of quantum wells and electron density [39]. We will see later that this clear agreement of the observed MP frequency with the expected $\omega_{B}$ is in contrast to the cavity case, where the cavity modes play a prominent role in reshaping the MP optical response.
@@ III THz Magnetospectroscopy of Strongly Coupled Cavity
Figure 3: Interaction of cavity modes with MP resonance through magnetic field tuning: a). Measured reflectance spectra at 3K as a function of magnetic field strength. Dashed line indicates the cyclotron frequency, $\omega_{B}$. TM (LP) indicates the lower polariton branch of the ISB polariton with the TM mode, TE indicates the TE mode. b) Calculated spectra using the formalism in Section IV.1, accounting for Coulombic interaction. For the theory parameters, see Table 1 in App. C. Both TM upper (UP) and lower (LP) polaritons are visible, as well as TE mode. c) Close-up of non-smoothed spectra of lower polariton branch. In (a,c), the black dashed line indicates the bare MP resonance as illustrated in Fig.2(c). d) Individual spectra of TM LP for three values of the magnetic field corresponding to $\omega_{B}<\omega_{TM}$ (purple), $\omega_{B}=\omega_{TM}$ (orange) and $\omega_{B}>\omega_{TM}$ (yellow).
@@
The cavity-coupled QW device is then measured by reflection-mode THz-TDS at 3K with an applied DC magnetic field from 0 – 9 T. In this magnetic field range, the MP resonance sweeps across the entire spectral range of the measurement and interacts with the different cavity modes consecutively. The resulting spectra are shown in Figure 3a. The MP resonance frequency, given by $\omega_{B}=eB/m^{*}$, is shown by the dashed line. For both the TE and TM modes, we observe a clear coupling of the cavity mode to the MP; however, the nature of the coupling is different depending on the mode.
@@
The high contrast feature observed in the spectra at 2.4 THz corresponds to the lower ISB polariton, where the TM cavity mode is coupled to the ISBT. The much lower contrast upper polariton is less visible in the spectra, and we therefore restrict our analysis to the lower polariton only. Although the TM mode is primarily polarized orthogonally to the MP, it still supports sufficient in-plane field components to couple to the MP. The observed coupling around 4 - 6 T (shown more clearly in Fig. 3c,d) therefore corresponds to a tripartite coupling between the ISBT and cavity (which form the ISB polariton) with the MP. Given its weak strength, this LP-MP coupling does not result in the typical anti-crossing shape, but rather in an enhanced broadening of the spectral cavity feature due to its mixing with the broadened MP resonance. Most interestingly, the location of this broadening feature is displaced from the crossing point with the bare MP frequency $\omega_{B}$, and can be understood as resulting from a significant blue-shift of the MP which is the signature of the non-local effects stemming from the Coulomb interaction. At larger B values, the interaction with the MP is still visible as a weak red-shift of the ISB polariton frequency, quantified by approximately 40 GHz.
@@
On the other hand, the higher-frequency TE mode at 3.5 THz is a purely photonic mode that has no contribution from the ISBT. It contains mostly in-plane fields which directly couple to the MP resonance, resulting in strong coupling and a clear anti-crossing behaviour. In contrast to the spectral region around the TM mode, here the MP appears to fit well the $\omega_{B}$ dependency on magnetic field. Fitting the response using a Hopfield model (assuming an MP frequency $\omega_{B}$) results in a polariton gap of 0.6 THz and therefore normalized coupling of $\eta=0.04$. We note that the value of the uncoupled cavity mode frequency extracted from the experiment in Fig. 3a is slightly shifted from the value 3.8 THz of ab-initio simulations in Fig. 1b. This is likely due to slightly mismatched cavity dimensions from their nominal values.
@@
The physical understanding of the different coupling of the TM and TE mode to the MP and ISBT resonances in the QW will be discussed in the next following Section IV.1.
@@ IV.1 Cavity-induced Coulomb interactions
From the theory of macroscopic dielectrics, it is well known that a polarizable medium is strongly affected by its own internal electric field. This self-interaction is due to the fact that the dipolar constituents of the material interact among each other via the Coulomb force, which is often approximated by the dipole-dipole interaction. When the medium is excited by a radiative source, it develops a macroscopic polarization density vector $\textbf{P}(\textbf{r},t)$, oscillating together with the external drive. This macroscopic motion of charges also gives rise to a macroscopic electric field inside the dielectric [40]:
@@
which splits into bulk and surface contributions.
@@
holding for $L_{c}k\ll 1$. The first term of Eq. (4) accounts for the free space contribution of Coulomb interactions, while the second term, proportional to $(L_{c}k)^{2}$, is the screening correction coming from the metallic boundaries. This phenomenology is also commonly known as non-locality [25].
@@
Figure 4: a) Cavity containing parabolic quantum well heterostructure: schematic (above) and simulated cavity fields (below) b) Experimentally measured spectra with magnetic field for cavity in a. Arrows indicate positions of line cuts in (c) for lower polariton (bottom) and upper polariton (top). c) Individual spectra of TM LP (left) and UP (right) for three values of the magnetic field corresponding to the purple, orange and yellow arrows in Fig. 4a. d) Calculated transmission of sample in b. For the theory parameters, see Table 1 in App. C.
@@ IV.2 Modeling the Coupled System
To demonstrate this, in Figure 3b we calculate the spectra taking into account the effect of Coulomb interaction and resulting wavelength-dependent MP frequency. This calculation is performed with a simple classical linear model, starting from Maxwell equations in the presence of a polarizable material and projecting them on the cavity two-mode subspace (see App. A). While the cavity modes are transverse modes [45], we do not discard the longitudinal component of the electric field, which gives rise to what we call the Coulomb contribution mentioned in Eq. (1). As indeed described in the previous section IV.1, and detailed in Appendix A-B, it is this longitudinal contribution which provides a $k$-dependent depolarization shift affecting the MP during the interaction with the TM mode.
@@
We see a good agreement of the measured spectra with the calculated spectra, in Fig. 3a and Fig. 3b, respectively. Comparing more closely in Fig. 3d, the calculated spectra accurately demonstrate the broadened linewidth of the ISB polariton as a result of the interaction with the Coulomb-shifted MP. The TE mode is also well reproduced, showing a clear anti-crossing at the MP frequency, in contrast to the TM mode. This demonstrates that to accurately model light-matter coupled systems with electronic transitions and inhomogeneous electric fields, it is crucial to include the influence of the field inhomogeneity on the Coulomb interactions in the 2DEG.
IV.3 Generality of Non-Locality

### v2

@@ I Introduction
In Landau polariton systems, the description of the coupling relies on Kohn’s theorem, which states that the cyclotron resonance frequency in a translationally-invariant two-dimensional gas is unaffected by electron-electron interactions [28]. Despite its simplicity, this result turns out to be extremely powerful, reducing the description of the system to a non-interacting one, and allowing for the development of a simple, intuitive picture of light-matter interactions in terms of a Hopfield-like polaritonic description [29, 30]. At the same time, Kohn’s theorem is also a limitation, excluding the possibility to observe non-linearities and possible many-body physics in the optical properties of these systems. Going beyond its range of validity is thus becoming an important topic of research [31], which turns out to be surprisingly challenging: besides the use of intense THz sources as in Ref. [32], the fundamental strategy to break Kohn’s theorem and observe the effect of Coulomb interactions in the cyclotron resonance is to design the system to explicitly break translational invariance. This can be naturally achieved in resonant cavities with highly sub-wavelength field confinement and inhomogeneous fields [33], where the violation of Kohn’s theorem is manifested as a non-local polaritonic response [25, 34, 35].
@@
Using THz time-domain spectroscopy, we observe a pronounced renormalization of the magnetoplasmon resonance frequency through electron–electron Coulomb interactions. This is activated due to the spatial inhomogeneity of the TM cavity mode, which is strongly hybridized with the ISB transition. Unlike Ref. [25], where non-locality resulted from a deeply sub-wavelength cavity miniaturization design ($\lambda/1000$), here the violation of Kohn’s theorem occurs with moderate field confinement ($\lambda/60$), provided that the resonant mode is sufficiently inhomogeneous. Furthermore, we demonstrate that by tuning the magnetoplasmon with the magnetic field, it is possible to strongly couple it with a different TE mode, which is mostly homogeneous and in-plane polarized. As a consequence, the TE mode is not hybridized with the ISBT, and thus exhibits purely cavity-like behavior. Moreover, the non-locality disappears and Kohn’s theorem is re-established due to a high level of spatial homogeneity.
@@
Figure 2: a) Schematic of experimental reflection THz-TDS set up. THz transmitter and receiver indicated with a ‘T’ and ‘R’ respectively. b) Measured reflectance spectra of the QW-loaded cavity as a function of temperature (with no applied magnetic field). The lower polariton branch (LP) is indicated. Below schematic indicating the frequency mismatch between the cavity and ISB resonances. Inset - single spectra slice at 10 K. c) Measured THz-TDS transmission spectra of semiconductor heterostructure in the absence of a cavity at 3K. The black dashed line indicates the unperturbed MP frequency given by $\omega_{B}=eB/m^{*}$, where $m^{*}=0.067m_{e}$. Inset - single spectra slice at 3T. Regions in (b) and (c) where reflectance and transmission are greater than 1.0 are unphysical and due to normalization to reference traces with low signal. However, these are included for clarity of the data.
@@ II.2 Quantum Well
[…]
@@
When a magnetic field is applied, the QW also exhibits an MP resonance due to the collective cyclotron resonance of the electrons [38]. Unlike the ISBT, which is fixed in frequency with the material growth, the MP resonance frequency increases linearly with applied magnetic field as $\omega_{B}=eB/m^{*}$, where $m^{*}=0.067m_{e}$ is the effective mass of electrons in the GaAs QW. The MP resonance in the bare QW heterostructure without the cavity is measured using transmission THz-TDS is shown in Fig. 2c. The THz response shows an almost perfect agreement with the expected linear dependence on the magnetic field (shown as a dashed black line). The linewidth can be estimated to have an almost $B$-independent value of $\kappa/(2\pi)\approx 300\,$GHz (see inset), which is the expected order of magnitude considering the number of quantum wells and electron density [39]. We will see later that this clear agreement of the observed MP frequency with the expected $\omega_{B}$ is in contrast to the cavity case, where the cavity modes play a prominent role in reshaping the MP optical response.
@@ III THz Magnetospectroscopy of Strongly Coupled Cavity
Figure 3: Interaction of cavity modes with MP resonance through magnetic field tuning: a). Measured reflectance spectra at 3K as a function of magnetic field strength. Dashed line indicates the cyclotron frequency, $\omega_{B}$. TM (LP) indicates the lower polariton branch of the ISB polariton with the TM mode, TE indicates the TE mode. b) Calculated spectra using the formalism in Section IV.1, accounting for Coulombic interaction. For the theory parameters, see Table 1 in App. C. Both TM upper (UP) and lower (LP) polaritons are visible, as well as TE mode. The white dashed line is the bare cyclotron $\omega_{B}$, while the red solid line is a guide for the eye, obtained by interpolating between the bare cyclotron frequency and the brightest, blue-shifted, cyclotron k-mode (see Appendix B for further details). c) Close-up of non-smoothed spectra of lower polariton branch. In (a,c), the black dashed line indicates the bare MP resonance as illustrated in Fig. 2(c). d) Individual spectra of TM LP for three values of the magnetic field corresponding to $\omega_{B}<\omega_{TM}$ (purple), $\omega_{B}=\omega_{TM}$ (red) and $\omega_{B}>\omega_{TM}$ (yellow).
@@
The cavity-coupled QW device is then measured by reflection-mode THz-TDS at 3K with an applied DC magnetic field from 0 – 9 T. In this magnetic field range, the MP resonance sweeps across the entire spectral range of the measurement and interacts with the different cavity modes for different values of the B field. The resulting spectra are shown in Figure 3a. The MP resonance frequency, given by $\omega_{B}=eB/m^{*}$, is shown by the dashed line. For both the TE and TM modes, we observe a clear coupling of the cavity mode to the MP; however, the nature of the coupling is different depending on the mode.
@@
The high contrast feature observed in the spectra at 2.4 THz corresponds to the lower ISB polariton, where the TM cavity mode is coupled to the ISBT. The much lower contrast upper polariton is less visible in the spectra, and we therefore restrict our analysis to the lower polariton only. Although the TM mode is primarily polarized orthogonally to the MP, it still supports sufficient in-plane field components to couple to the MP. The observed coupling around 4 - 6 T (shown more clearly in Fig. 3c,d) therefore corresponds to a tripartite coupling between the ISBT and cavity (which form the ISB polariton) with the MP. This LP-MP coupling does not result in the typical anti-crossing shape, but rather in an enhanced broadening of the spectral cavity feature due to its mixing with the broadened MP resonance. Most interestingly, the location of this broadening feature is displaced from the crossing point with the bare MP frequency $\omega_{B}$, as seen most clearly in the interpolated color map of Fig. 3a. This can be understood as resulting from a significant blue-shift of the MP which is the signature of the non-local effects stemming from the Coulomb interaction, as highlighted by the red line in the simulation panel in Fig. 3b. The origin of this blue-shift will be discussed in more detail in Section IV.
@@
On the other hand, the higher-frequency TE mode at 3.5 THz is a purely photonic mode that has no contribution from the ISBT. It contains mostly in-plane fields which directly couple to the MP resonance, resulting in strong coupling and a clear anti-crossing behaviour. In contrast to the spectral region around the TM mode, here the MP appears to fit well the $\omega_{B}$ dependency on magnetic field. Fitting the response using a Hopfield model (assuming an MP frequency $\omega_{B}$) results in a polariton gap of 0.6 THz and therefore normalized coupling of $\eta=0.04$.
@@
The physical understanding of the different coupling of the TM and TE mode to the MP and ISBT resonances in the QW will be discussed in the following Section IV.
@@ IV.1 Cavity-induced Coulomb interactions
From the theory of macroscopic dielectrics, it is well known that a polarizable medium is strongly affected by its own internal electric field. This self-interaction is due to the fact that the dipolar constituents of the material interact among each other via the Coulomb force, which is often approximated by the dipole-dipole interaction. When the medium is excited by a radiative source, it develops a macroscopic polarization density vector $\textbf{P}(\textbf{r},t)$, oscillating together with the external drive. This macroscopic motion of charges also gives rise to a macroscopic electric field inside the dielectric, which splits into bulk and surface contributions [40]:
@@
holding for $L_{c}k\ll 1$. The first term of Eq. (4) accounts for the free space contribution of Coulomb interactions, while the second term, proportional to $(L_{c}k)^{2}$, is the screening correction coming from the metallic boundaries (not to be confused with the surface contribution in Eq. (1), which refers only to the QW dielectric properties, while the metallic boundary affects the $(\textbf{r}-\textbf{r}^{\prime})/|\textbf{r}-\textbf{r}^{\prime}|^{3}$ kernel in Eq. (1), being replaced with the proper Green’s function [see Appendix A.2]). This phenomenology is also commonly known as non-locality [25].
@@
Figure 4: a) Cavity containing parabolic quantum well heterostructure: schematic (above) and simulated cavity fields (below) b) Experimentally measured spectra with magnetic field for cavity in a. Arrows indicate positions of line cuts in (c) for lower polariton (bottom) and upper polariton (top). c) Individual spectra of TM LP (left) and UP (right) for three values of the magnetic field corresponding to the purple, red and yellow arrows in Fig. 4a. d) Calculated transmission of sample in b. For the theory parameters, see Table 1 in App. C.
@@ IV.2 Modeling the Coupled System
To demonstrate this, in Figure 3b we calculate the spectra taking into account the effect of Coulomb interaction and resulting wavelength-dependent MP frequency. This calculation is performed with a simple classical linear model, starting from Maxwell equations in the presence of a polarizable material and projecting them on the cavity two-mode subspace (see App. A). The transmission/reflection is then extracted following Appendix B by solving the following coupled-modes equations
$\begin{split}\left(\partial_{t}^{2}+\omega_{\lambda}^{2}\right)\tilde{\mathcal{E}}_{\lambda}&=-\partial_{t}^{2}\sum_{k}\left(\tilde{w}_{\lambda,k}\tilde{P}_{k}^{\rm mp}+\tilde{u}_{\lambda,k}\tilde{P}_{k}^{\rm isb}\right),\\ \left(\partial_{t}^{2}+\bar{\omega}_{B}^{2}(k)\right)\tilde{P}_{k}^{\rm mp}&=N_{\rm qw}\omega_{P}^{2}\sum_{\lambda}\tilde{w}_{\lambda,k}\tilde{\mathcal{E}}_{\lambda},\\ \left(\partial_{t}^{2}+\omega_{\rm isb}^{2}\right)\tilde{P}_{k}^{\rm isb}&=N_{\rm qw}\omega_{P}^{2}\sum_{\lambda}\tilde{u}_{\lambda,k}\tilde{\mathcal{E}}_{\lambda}.\end{split}$ (5)
Here $\tilde{\mathcal{E}}_{\lambda}$ is the electric field amplitude of the $\lambda$-th cavity mode, with frequency $\omega_{\lambda}$; $\tilde{P}_{k}^{\rm mp}$ is the dielectric polarization amplitude of the MP $k$-th mode, with frequency $\bar{\omega}_{B}(k)$ which includes the $k$-dependent non-local Coulomb contribution; $\tilde{P}_{k}^{\rm isb}$ is the dielectric polarization amplitude of the ISB transition with frequency $\omega_{\rm isb}$. As detailed in the Appendix A.4 $\tilde{w}_{\lambda,k},\tilde{u}_{\lambda,k}$ are coefficients that depend on the overlap between the $\lambda$-th spatial cavity mode and the $k$-th MP/ISB mode. These coefficients carry the fundamental information about the cavity-induced spatial inhomogeneity which is eventually able to trigger the observed non-locality.
It is worth stressing that, while the cavity modes are transverse modes [45] (whose dynamics are in the first line of Eq. (5)), we do not discard the longitudinal component of the electric field, which gives rise to what we call the Coulomb contribution mentioned in Eq. (1). As indeed described in the previous section IV.1, and detailed in Appendix A-B, it is this longitudinal contribution which provides a $k$-dependent depolarization shift affecting the MP frequency, $\bar{\omega}_{B}(k)$, and emerges during the interaction with the TM mode.
We see a good agreement of the measured spectra in Fig. 3a with the calculated spectra in Fig. 3b. Comparing more closely in Fig. 3d, the calculated spectra accurately demonstrate interaction of the ISB polariton with the Coulomb-shifted and broadened MP. In addition, the calculation also models the blue-shift of the ISB polariton at low magnetic field, when the MP frequency is much smaller than the ISB one and $\omega_{B}\rightarrow 0$. This shift is compatible with the standard polariton gap due to ultrastrong coupling, and is present even though we don’t observe clear polariton splitting due to the broadened cyclotron linewidth. However, since this shift is much smaller than the linewidth of the ISB polariton itself and on the order of our experimental system’s spectral resolution, we cannot experimentally determine its quantitative value with accuracy.
The TE mode is also well reproduced, showing a clear anti-crossing at the MP frequency, in contrast to the TM mode. We note for the calculation in 3b we use a value of 3.508 THz for the uncoupled TE cavity mode frequency, which is extracted from the experiment in Fig. 3a. This is slightly shifted from the value 3.8 THz of ab-initio simulations in Fig. 1b, likely due to slightly mismatched cavity dimensions from their nominal values. This and other simulation parameters used can be found in Table 1 in Appendix C.
@@
For comparison, in Appendix D we also show the calculated spectra without including the Coulomb interaction, which reproduces the TE spectra, but does not accurately reproduce the experimental spectra for the TM mode. This demonstrates that to accurately model light-matter coupled systems with electronic transitions and inhomogeneous electric fields, it is crucial to include the influence of the field inhomogeneity on the Coulomb interactions in the 2DEG.
IV.3 Hamiltonian Description and Ultra-Strong Coupling
From the classical dielectric theory in Eq. (5) (see in Appendix A for more details), we can also derive the respective Hamiltonian for the quantized system, which allows us to understand the dominant energies in the system. Introducing the annihilation operators $a_{\lambda},b_{k},c_{k}$, respectively for the $\lambda$-th cavity mode, MP and ISB $k$-th mode, the total system’s Hamiltonian is given as the sum of four contributions $H=H_{c}+H_{\rm mp}+H_{\rm isb}+H_{\rm I}$, where (see Appendix E for the derivation)
$\begin{split}H_{c}&=\sum_{\lambda}\hbar\omega_{\lambda}a_{\lambda}^{{\dagger}}a_{\lambda}\\ H_{\rm mp}&=\sum_{k}\hbar\omega_{B}b^{{\dagger}}_{k}b_{k}\\ H_{\rm isb}&=\sum_{k}\hbar\omega_{\rm isb}c^{{\dagger}}_{k}c_{k},\end{split}$ (6)
are, respectively, the cavity, MP, and ISB individual Hamiltonians. The light-matter interaction Hamiltonian, describing the coupling between the electrodynamical cavity field and the MP, ISB excitations, is given by
$\begin{split}&H_{\rm I}=\\ &-i\frac{\hbar\tilde{\omega}_{P}}{2}\sum_{\lambda,k}\left(a_{\lambda}-a_{\lambda}^{{\dagger}}\right)\left[\tilde{w}_{\lambda,k}\left(b_{k}+b_{k}^{{\dagger}}\right)+\tilde{u}_{\lambda,k}\left(c_{k}+c_{k}^{{\dagger}}\right)\right]\\ &+\sum_{\lambda}\frac{\hbar\tilde{\omega}_{P}^{2}}{4\omega_{\lambda}}\left[\sum_{k}\left(\tilde{w}_{\lambda,k}\left(b_{k}+b_{k}^{{\dagger}}\right)+\tilde{u}_{\lambda,k}\left(c_{k}+c_{k}^{{\dagger}}\right)\right)\right]^{2}\\ &+\frac{\hbar\tilde{\omega}_{P}^{2}}{4\omega_{B}}\sum_{k}\zeta_{k}\left(b_{k}+b_{k}^{{\dagger}}\right)^{2}.\end{split}$ (7)
Here we introduced the multiple-QW collective plasma frequency as $\tilde{\omega}_{P}=\sqrt{N_{\rm qw}}\omega_{P}$.
From Eq. (7) we see that the light-matter interaction is not given by a single term, but it is rather structured. In the first line, we have the direct linear coupling between the cavity electric field and the total matter’s polarization, obtained as a weighted sum of the MP and ISB polarizations:
$\begin{split}H_{\rm lin}=&-i\frac{\hbar\tilde{\omega}_{P}}{2}\sum_{\lambda,k}\left(a_{\lambda}-a_{\lambda}^{{\dagger}}\right)\\ &\times\left[\tilde{w}_{\lambda,k}\left(b_{k}+b_{k}^{{\dagger}}\right)+\tilde{u}_{\lambda,k}\left(c_{k}+c_{k}^{{\dagger}}\right)\right].\end{split}$ (8)
This term scales linearly with the plasma frequency (the light-matter coupling constant), $\omega_{P}$, and, under the rotating-wave approximation (RWA) gives the photon-matter flip-flop interaction that defines the polariton excitation [46].
In the second line of Eq. (7), we have a matter-only term, scaling quadratically with the collective plasma frequency, $\tilde{\omega}_{P}^{2}$
$H_{P^{2}}=\sum_{\lambda}\frac{\hbar\tilde{\omega}_{P}^{2}}{4\omega_{\lambda}}\left[\sum_{k}\left(\tilde{w}_{\lambda,k}\left(b_{k}+b_{k}^{{\dagger}}\right)+\tilde{u}_{\lambda,k}\left(c_{k}+c_{k}^{{\dagger}}\right)\right)\right]^{2}.$ (9)
This term is the famous $P^{2}$-term (polarization square), ensuring the stability of the system, and that also rules out any possible superradiant phase transition in the ground-state [6, 41, 43, 47].
The last term in Eq. (7) describes the Coulomb non-locality emerging from Eqs. (1)-(2)-(4)
$H_{\rm nloc.}=\frac{\hbar\tilde{\omega}_{P}^{2}}{4\omega_{B}}\sum_{k}\zeta_{k}\left(b_{k}+b_{k}^{{\dagger}}\right)^{2},$ (10)
The non-locality term is quadratic in the collective plasma frequency, $\tilde{\omega}_{P}^{2}$, and it can also be labelled as a $P^{2}$-term. However, contrary to the $P^{2}$-term in Eq. (9), which emerges from the transverse part of the electromagnetic interactions [45, 43, 48], the non-locality emerges from the longitudinal part of the electric field [48], and thus, while similar, these two terms are very different in nature and must not be confused.
The $\tilde{\omega}_{P}^{2}$, quadratic scaling of the non-locality term in Eq. (10), clearly links its phenomenology with the so-called ultra-strong coupling (USC) regime [2, 1, 5]. In this regime, the light-matter coupling strength becomes as large as the intrinsic cavity or QW energy scales (in our case, see Table 1, $\hbar\tilde{\omega}_{P}\sim 1\,{\rm THz}\sim\hbar\omega_{\lambda},\hbar\omega_{B},\hbar\omega_{\rm isb}$). In this way, the linear cavity-QW coupling given by Eq. (8) becomes sub-leading, indicating that, whenever the inhomogeneity of the system allows the cavity to excite $k>0$ modes (i.e. the $\tilde{w}_{\lambda,k},\tilde{u}_{\lambda,k}$ overlap coefficients are non-zero at finite $k$), the USC regime is dominated by non-locality effects, affecting non-perturbatively the phenomenology of cavity QED light-matter interactions [43].
IV.4 Generality of Non-Locality